# Flow-Aware Ellipsoidal Filtration for Persistent Homology of Recurrent Signals

### v1

@@ 4.1 Local Covariance Estimation
This captures recurrent states that reappear near the same region of the attractor at possibly very different times.
@@
which unites local temporal continuity with spatial recurrence. This union is crucial because it balances short-term temporal information with the broader geometric structure of the attractor.
@@
This covariance matrix encodes both the direction and magnitude of local variation, effectively capturing how the trajectory is oriented and stretched in phase space. The eigenvectors of $\Sigma_{i}$ indicate the dominant flow directions, while the eigenvalues reflect how rapidly the signal varies along each direction.
4.2 Adaptive Ellipsoids from Local Covariance
@@
4.3 Intersection of Adaptive Ellipsoids
@@ 4.4 Ellipsoid Complex
4.5 Algorithm for Building the Ellipsoid Complex
Algorithm 1 summarises the computational procedure for building the ellipsoid complex at a given scale. For each pair of points, the algorithm checks whether their adaptive ellipsoids intersect according to the convex criterion in Eq. (11). The resulting edge list defines the 1-skeleton, from which the flag complex is constructed. The flag complex construction and persistent homology were implemented with the SimplexTree data structure from the GUDHI library (Maria et al., 2014).
@@ 4.5 Algorithm for Building the Ellipsoid Complex
Repeating this process over an increasing sequence of scales $0<\varepsilon_{1}<\varepsilon_{2}<\cdots$ produces the filtration $\{\mathcal{K}_{\varepsilon_{t}}\}$ used for persistent homology computation. This filtration captures how topological features of the data—such as loops and connected components—emerge and disappear as neighbourhoods expand along the flow’s geometry.
Figure 2: Ellipsoids fitted to local neighbourhoods. Each neighbourhood includes the temporal window of three past and three future points, together with the 15 nearest spatial neighbours. The ellipsoids align with the local geometry estimated from these combined spatio–temporal neighbourhoods.
Code availability. The implementation and analysis scripts are (to be) hosted at github.com/omerberyilmaz/flow-aware-ellipsoidal-filtration.
5 Experiments and Results
5.1 Filtration behaviour with dense and sparse phase–space regions
We generate a two-dimensional trajectory from a Hamiltonian system with a non-linear potential, discretised using the Störmer–Verlet method:
@@ 5.1 Filtration behaviour with dense and sparse phase–space regions
with parameters $T=17.22$, $h=0.08$, $r=5$, $\epsilon=0.39$, $\omega=3$, and initial conditions $(q_{0},p_{0})=(0.5,1.0)$. Gaussian noise is added independently to both coordinates to mimic measurement uncertainty. The resulting trajectory is shown in phase space $(p,q)$ in Figure 3.
Although the trajectory is sampled at uniform time intervals, the induced point cloud in phase space is highly non-uniform: slow dynamics yield dense regions, whereas fast transitions leave sparse ones. Because temporal information is ignored, a purely spatial approach cannot distinguish whether nearby points belong to the same dynamical phase. This limitation makes the system a useful test case for evaluating how different filtrations recover the trajectory’s topology from its point cloud representation. Importantly, the same fast–slow structure also arises in practice, for example in 3D accelerometer recordings of cyclic human motion.
@@
Figure 3 compares the three filtrations across increasing neighbourhood scales. The aim is to recover the global loop that reflects the system’s recurrence, despite the narrow bottleneck in the middle. With Fermat distance (Figure 3a), dense regions connect quickly, but the bottleneck is bridged before sparse regions are included, preventing loop recovery. Vietoris–Rips (Figure 3b) also connects across the bottleneck too early and likewise fails to recover the loop. In contrast, the ellipsoidal filtration (Figure 3c) adapts neighbourhoods to local flow direction and speed. This allows the loop to appear at smaller scales while the bottleneck remains separated in the leftmost plot, giving a more accurate representation of the underlying topology.
These results show that neighbourhoods defined only by density or distance cannot recover meaningful topology, because the scale parameter does not reflect the local flow structure. Ellipsoidal filtration combines temporal information with spatial proximity, enabling the correct loop structure to be recovered even in the presence of bottlenecks. Although this dataset is synthetic, similar structures are common in real-world signals, such as accelerometer data from running and cycling (see Section 5.3). The following experiments build on this result, showing how appropriate neighbourhood selection is crucial for denoising recurrent signals and estimating recurrence times.
(a) Fermat
(b) Vietoris–Rips
(c) Ellipsoidal
Figure 3: Comparison of three filtrations on the same dataset: (a) Fermat distance, (b) Vietoris–Rips, and (c) Ellipsoidal filtration. Each panel shows a $1\times 4$ grid across increasing scale values.
5.2 Denoising a recurrent signal
@@
The synthetic signal is designed to expose two representative challenges: a linearly increasing frequency and a narrow bottleneck in the trajectory. We define
$x(t)=10\cos\!\big(\phi(t)\big),\qquad y(t)=2\sin\!\big(\phi(t)\big)\,S\!\big(x(t)\big),\qquad t\in[0,2],$ (16)
with $n=500$ uniform samples. The instantaneous frequency is
$f(t)=f_{\mathrm{start}}+\frac{f_{\mathrm{end}}-f_{\mathrm{start}}}{t_{\max}}\,t,\quad(f_{\mathrm{start}},f_{\mathrm{end}})=(1,10)\ \text{Hz},\quad t_{\max}=2,$ (17)
and the chirp phase is
@@
$S(x)=1-0.9\exp\!\left(-\tfrac{1}{2}\left(\frac{x}{\max_{t}|x(t)|}\right)^{2}\right).$ (19)
@@
We compare five strategies: (i) a Vietoris–Rips-based (isotropic, “spherical”) topological filter, (ii) an ellipsoidal (flow-aware) topological filter built using the spatio–temporal PCA construction as described in Section 4, (iii) a $k$-NN topological filter with $k=20$, (iv) a fixed 20-sample moving average, and (v) an adaptive moving average whose window length is set from an FFT-based frequency estimate (Nyquist rule). For Vietoris–Rips and ellipsoidal topological filters, the neighbourhood scale is chosen by the death time of the most persistent $H_{1}$ class.
Even when the noise level is modest (20 dB), the methods differ markedly in how well they preserve the input structure (Figure 4). Panel (a) shows the noisy input with a clear frequency rise and a narrow low-amplitude bottleneck.
@@
In panel (b), the fixed 20-sample moving average is the most clearly destructive: once the period drops below the window length, the filter averages across opposite lobes and the waveform is driven towards zero in the middle segment; fine peak details disappear. In panel (c), the adaptive moving average is not adaptive enough to preserve the signal structure. As the frequency increases, it fails to adjust its window size accordingly, leading to reduced peak-to-peak amplitude, rounded peaks, and loss of fine details.
@@
Topological filters behave differently for the low-amplitude component (see Fig. 4(b)). The $k$-NN filter in panel (d) preserves oscillations over large portions of the record, yet in the bottleneck its fixed-size neighbourhood spans opposing directions of the trajectory, causing local attenuation and shape distortion. The VR (spherical) filter in panel (e) is more severely affected because isotropic balls cut across the thin geometry; cross-bottleneck averaging reduces peak-to-peak amplitude in the low-amplitude region. By contrast, the ellipsoidal filter in panel (f) aligns neighbourhoods with the local flow direction and speed, avoiding cross-bottleneck mixing; amplitude and phase are preserved across the record, including the bottleneck.
@@
Figure 5 reports root-mean-square error (RMSE) versus SNR for $x(t)$ and $y(t)$. For $x(t)$, the moving-average baselines can yield lower RMSE at low SNRs because aggressive smoothing suppresses noise; however, their error quickly plateaus as SNR increases. The three topological filters improve steadily with SNR and overtake the baselines from moderate SNRs upwards, with broadly similar performance among the topological methods on $x(t)$. For $y(t)$, the contrast is stronger: the ellipsoidal filter achieves the lowest RMSE over the mid- to high-SNR range, while $k$-NN and spherical VR also surpass the time-domain baselines once the SNR is moderate. This behaviour reflects a trade-off: time-domain smoothing reduces noise but also attenuates peak-to-peak structure, so its RMSE gains saturate, whereas topology-based neighbourhoods preserve structure and therefore continue to improve as noise decreases.
(a) Filtered outputs for the high-amplitude component $x(t)$.
(b) Filtered outputs for the high-amplitude component $y(t)$.
@@
Figure 4: Comparison of filters applied to denoise the signal $x(t)$ and $y(t)$ at SNR = 20 dB. (a) Original and noisy signals, (b) Moving average,(c) Adaptive moving average, (d) $k$-NN topological filter, (e) Spherical topological filter, (f) Ellipsoidal topological filter.
(a) RMSE vs. SNR for the high-amplitude component $x(t)$.
(b) RMSE vs. SNR for the low-amplitude component $y(t)$.
Figure 5: Quantitative comparison of filters across SNR levels. Ellipsoidal filtration achieves the lowest RMSE, particularly for the low-amplitude component $y(t)$.
5.3 3D accelerometer data denoising
@@
We now evaluate the method on real data: publicly available tri-axial accelerometer recordings from cycling (Vieten and Weich, 2019). The data were sampled at 500 Hz; we analysed a 3,500-sample segment from the first participant.
To set comparable neighbourhood scales for Vietoris–Rips (VR) and the ellipsoidal filtration, we first computed persistent homology on a shorter 350-sample segment (about one full cycle). Let $B$ and $D$ be the birth and death of the dominant $H_{1}$ class, and L=DBL=D-B its lifetime. For each filtration we examined four scales: $B$, $B+\tfrac{1}{2}L$, $D$, and $D+\tfrac{1}{2}L$. When estimating ellipsoidal covariances, we pooled neighbours across adjacent cycles to stabilise local variance estimates, as described in the Methods.
@@
Figure 6 shows persistence diagrams for the same 3D segment under the Vietoris-Rips and the proposed Ellipsoidal filtrations (red: $H_{0}$, blue: $H_{1}$). In the VR diagram, two $H_{1}$ classes are clearly more persistent, whereas in the ellipsoidal diagram a single $H_{1}$ class dominates. The blue points farthest from the diagonal represent the most persistent loop in each panel. In the denoising and recurrence time estimation 5.4 we consider the scale of the most persistent $H_{1}$ class.
@@
Figure 7 compares spherical and ellipsoidal neighbourhood averaging across four scales. At small scales, both methods preserve the data’s loop structure. As the scale increases, spherical neighbourhoods span opposite sides of narrow passages and average across bottlenecks, compressing the geometry and attenuating amplitude. In contrast, ellipsoidal averaging follows the local flow and better preserves loops across scales.
@@
Figure 8 applies denoising by taking the geometric median of the samples within each sphere or ellipsoid. The geometric median yields smoother reconstructions and exhibits greater robustness to noise than the mean.
(a) VR filtration persistence diagram
(b) Ellipsoidal filtration persistence diagram
Figure 6: Persistence diagrams computed on the 3D accelerometer signal. (a) Vietoris–Rips (spherical) filtration. (b) Ellipsoidal filtration.
(a) s=2.494,e=2.052s=2.494,\ e=2.052
(b) s=3.626,e=3.334s=3.626,\ e=3.334
(c) s=4.759,e=4.616s=4.759,\ e=4.616
(d) s=5.891,e=5.898s=5.891,\ e=5.898
Figure 7: Accelerometer denoising using mean averaging: comparison of spherical versus ellipsoidal neighbourhoods at increasing scales. For each filtration, the scale is set relative to the most persistent $H_{1}$ feature: $b$, $\tfrac{1}{2}\ell, b+b+\ell$, and $b+\tfrac{3}{2}\ell$, where $b$ is the birth scale and $\ell$ the lifetime. The corresponding numerical values are shown in subcaptions. The left column shows the raw data, the middle column the spherical neighbourhood outcomes, and the right column the ellipsoidal neighbourhood outcomes.
(a) s=2.494,e=1.967s=2.494,\ e=1.967
(b) s=3.626,e=3.419s=3.626,\ e=3.419
(c) s=4.759,e=4.872s=4.759,\ e=4.872
(d) s=5.891,e=6.325s=5.891,\ e=6.325
Figure 8: Accelerometer denoising using geometric median averaging: comparison of spherical versus ellipsoidal neighbourhoods at increasing scales. For each filtration, the scale is set relative to the most persistent $H_{1}$ feature: $b$, $\tfrac{1}{2}\ell, b+b+\ell$, and $b+\tfrac{3}{2}\ell$, where $b$ is the birth scale and $\ell$ the lifetime. The corresponding numerical values are shown in subcaptions. The left column shows the raw data, the middle column the spherical neighbourhood outcomes, and the right column the ellipsoidal neighbourhood outcomes.
Figure 9: Accelerometer time-series at scale corresponding to the most persistent $H_{1}$ feature (s=5.891,e=5.898s=5.891,\ e=5.898). The three panels show the $x$-, $y$-, and $z$-axis accelerations respectively.
5.4 Recurrence-time estimation
@@
We assess whether the same flow-aware neighbourhoods that improved denoising (Sections 5.25.3) also improve recurrence-time estimation. The concept of recurrence and first-return times was introduced in Section 2.1. For consistency, we use the chirp signal defined in Section 5.2, which combines a linearly increasing frequency, anisotropic deformation, and a narrow bottleneck as described earlier.
@@
$T_{1}(i)=\min\{\,j>i:j-i\geq\tau_{\min},\ x_{j}\in\mathcal{N}_{i},\ \text{and}\ x_{k}\notin\mathcal{N}_{i}\ \ \forall\,i<k<j\,\},$ (22)
where $\tau_{\min}=15$ samples prevents spurious early returns introduced by noise or oversampling. For reference, ground-truth first returns are computed from the unwrapped phase $\theta(t)$ as the smallest $j>i$ such that $\theta[j]\geq\theta[i]+2\pi$. As in Section 5.3, we evaluate four scales per filtration taken from the dominant $H_{1}$ class: $B$, $B+\tfrac{1}{2}(D-B)$, $D$, and $D+\tfrac{1}{2}(D-B)$.
@@
Figures 11 and 12 show the recurrence structure across scales. On the clean signal, ellipsoidal neighbourhoods detect almost all first returns across scales except at the birth scale. In contrast, fixed-radius balls struggle as the period shortens: first returns are frequently missed at all scales, and at larger scales balls span opposite sides of the bottleneck, producing spurious early returns. Under noise, ellipsoidal neighbourhoods capture more first returns, missing only one first return at the largest scale. Importantly, no spurious early returns are caused by the bottleneck which makes it possible to estimate returns for all points.
(a) Clean — Vietoris–Rips
(b) Clean — Ellipsoidal
(c) Noisy — Vietoris–Rips
(d) Noisy — Ellipsoidal
Figure 10: Persistent diagrams for clean and noisy signals under two filtrations. Columns compare Vietoris–Rips (left) to ellipsoidal (right); rows compare clean (top) to noisy (bottom).
(a) $\varepsilon=0.223$
(b) $\varepsilon=0.519$
(c) $\varepsilon=0.815$
(d) $\varepsilon=1.112$
(e) $r=0.240$
(f) $r=0.305$
(g) $r=0.371$
(h) $r=0.437$
Figure 11: Clean signal: recurrence structure under ellipsoidal filtration (top, increasing $\varepsilon$ left to right) and Vietoris–Rips filtration (bottom, increasing $r$ left to right).
(a) $\varepsilon=0.223$
(b) $\varepsilon=0.445$
(c) $\varepsilon=0.667$
(d) $\varepsilon=1.112$
(e) $r=0.210$
(f) $r=0.284$
(g) $r=0.359$
(h) $r=0.434$
Figure 12: Noisy signal: recurrence structure under ellipsoidal filtration (top, increasing $\varepsilon$ left to right) and Vietoris–Rips filtration (bottom, increasing $r$ left to right).
6 Discussion

### v2

@@ IV.1 Local Covariance Estimation
This captures recurrent states that return to the same region of the attractor, even when they occur far apart in time.
@@
which unites local temporal continuity with spatial recurrence.
Spatial neighbours capture the variance arising from close returns, while temporal neighbours encode the local direction and rate of change. As the trajectory typically revisits the same regions many times, we choose $|\mathcal{S}_{i}|>|\mathcal{T}_{i}|$. Accordingly, the parameters $k$ and $\tau$ are selected to promote stable local tangent estimates, taking into account the sampling density and noise level of the data.
@@
This covariance matrix encodes both the direction and magnitude of local variation, effectively capturing how the trajectory is oriented and stretched in state-space. The eigenvectors of $\Sigma_{i}$ indicate the dominant flow directions, while the eigenvalues reflect how rapidly the signal varies along each direction.
IV.2 Adaptive Ellipsoids from Local Covariance
@@
IV.3 Intersection of Adaptive Ellipsoids
@@ IV.4 Ellipsoid Complex
IV.5 Algorithm for Building the Ellipsoid Complex
[…]
@@ IV.5 Algorithm for Building the Ellipsoid Complex
Repeating this process over an increasing sequence of scales $0<\varepsilon_{1}<\varepsilon_{2}<\cdots$ produces the filtration $\{\mathcal{K}_{\varepsilon_{t}}\}$ used for persistent homology computation. This filtration captures how topological features of the data—such as loops and connected components—emerge and disappear as neighbourhoods expand along the flow’s geometry. Fig. 2 illustrates this construction: the left panel shows ellipsoids placed along a recurrent trajectory at a fixed scale, while the right panel zooms into a single centre point and displays its temporal, spatial, and spatio–temporal neighbours. The ellipsoid at the centre point is obtained by applying PCA to the union of the temporal and spatial neighbour sets.
Figure 2: Flow-aware ellipsoids constructed from combined spatial and temporal neighbourhoods. Left: recurrent trajectory in the $(p,q)$-plane with ellipsoids placed along the orbit at a fixed scale. Right: zoomed view of the centre point, together with its temporal neighbours, spatial neighbours, and spatio–temporal neighbours (points that are close both in space and in time), as indicated in the legend. The ellipsoid is fitted by PCA to the union of the temporal and spatial neighbour sets, so that all ellipsoids share the same scale parameter but adapt their size and orientation to the local flow geometry.
Code availability. The implementation and analysis scripts will be released in a public repository after the paper is accepted.
V Experiments and Results
V.1 Filtration behaviour with dense and sparse phase–space regions
We generate a two-dimensional trajectory from a Hamiltonian system with a non-linear potential,
@@ V.1 Filtration behaviour with dense and sparse phase–space regions
with parameters $r=5$, $\epsilon=0.39$, and $\omega=3$. The system is discretised using the Störmer–Verlet method with time step $h=0.08$ over a total integration time $T=37.22$, and initial conditions $(q_{0},p_{0})=(0.5,1.0)$. Gaussian noise is added independently to both coordinates to mimic measurement uncertainty. The resulting trajectory, shown in state space $(p,q)$ in Fig. 3, is used to test the filtrations. The same dynamical system and numerical discretisation are used for the method demonstration in Fig. 2, with a slightly shorter integration time $T=27.22$.
Although the trajectory is sampled at uniform time intervals, the induced point cloud in state space is highly non-uniform: slow dynamics yield dense regions, whereas fast dynamics leave sparse ones. When temporal information is not incorporated, a purely spatial filtration may connect nearby points that belong to different phases of the trajectory. This tendency to connect geometrically close but dynamically distinct states makes the system a useful test case for evaluating how different filtrations recover the trajectory’s topology from its point cloud representation. Importantly, the same fast–slow structure also arises in practice, for example in 3D accelerometer recordings of cyclic human motion.
@@
Fig. 3 compares the three filtrations across increasing neighbourhood scales. The aim is to recover the global loop that reflects the system’s recurrence, despite the narrow bottleneck in the middle. With Fermat distance (Fig. 33(a)), dense regions connect quickly, but the bottleneck is bridged before sparse regions are included, preventing loop recovery. Vietoris–Rips (Fig. 33(b)) also connects across the bottleneck too early and likewise fails to recover the loop. In contrast, the ellipsoidal filtration (Fig. 33(c)) adapts neighbourhoods to local flow direction and speed. This allows the loop to appear at smaller scales while the bottleneck remains separated in the leftmost plot, giving a more accurate representation of the underlying topology.
In this dataset, neighbourhoods based only on density or Euclidean distance did not recover the target loop, because the scale parameter does not reflect the local flow structure. Ellipsoidal filtration combines temporal information with spatial proximity, enabling the correct loop structure to be recovered even in the presence of bottlenecks. Although this dataset is synthetic, similar structures are common in real-world signals, such as accelerometer data from running and cycling (see Section V.3). The following experiments build on this result, showing how appropriate neighbourhood selection is crucial for denoising recurrent signals and estimating recurrence times.
[…]
V.2 Denoising a recurrent signal
@@
The synthetic signal is designed to expose two representative challenges: a linearly increasing frequency and a narrow bottleneck in the trajectory. We define {subequations} {align} x(t) = 10cos(ϕ(t)),
y(t) = 2sin(ϕ(t)) S(x(t)), for $t\in[0,2]$ with $n=500$ uniform samples. The instantaneous frequency is
$f(t)=f_{\mathrm{start}}+\frac{f_{\mathrm{end}}-f_{\mathrm{start}}}{t_{\max}}\,t.$ (16)
Here $(f_{\mathrm{start}},f_{\mathrm{end}})=(1,10)\,\mathrm{Hz}$. The corresponding chirp phase is
@@
[…]
@@
We compare five strategies: (i) a Vietoris–Rips-based (isotropic, “spherical”) topological filter, (ii) an ellipsoidal (flow-aware) topological filter built using the spatio–temporal PCA construction as described in Section IV, (iii) a $k$-NN topological filter with $k=20$, (iv) a fixed 20-sample moving average, and (v) an adaptive moving average whose window length is set from an FFT-based frequency estimate (Nyquist rule). For Vietoris–Rips and ellipsoidal topological filters, the neighbourhood scale is chosen by the death time of the most persistent $H_{1}$ class.
@@
Even when the noise level is modest (20 dB), the methods differ markedly in how well they preserve the input structure (Fig. 4). The noisy input exhibits a clear frequency rise and a narrow low-amplitude bottleneck. The fixed 20-sample moving average is the most clearly destructive: once the period drops below the window length, the filter averages across opposite phases of the oscillation and the waveform is driven towards zero in the middle segment; fine peak details disappear. The adaptive moving average is not sufficiently responsive to preserve the signal structure. As the frequency increases, it fails to adjust its window size accordingly, leading to reduced peak-to-peak amplitude, rounded peaks, and loss of fine details.
@@
Topological filters behave differently for the low-amplitude component shown in the right panel of Fig. 4. The $k$-NN filter preserves oscillations over large portions of the record, yet in the bottleneck its fixed-size neighbourhood spans opposing directions of the trajectory, causing local attenuation and shape distortion. The VR (spherical) filter is more severely affected because isotropic balls cut across the thin geometry; cross-bottleneck averaging reduces peak-to-peak amplitude in the low-amplitude region. By contrast, the ellipsoidal filter aligns neighbourhoods with the local flow direction and speed, avoiding cross-bottleneck mixing; amplitude and phase are preserved across the record, including the bottleneck.
@@
Fig. 5 reports root-mean-square error (RMSE) versus SNR for $x(t)$ and $y(t)$. For $x(t)$, the moving-average baselines can yield lower RMSE at low SNRs because aggressive smoothing suppresses noise; however, their error quickly plateaus as SNR increases. The three topological filters improve steadily with SNR and overtake the baselines from moderate SNRs upwards, with broadly similar performance among the topological methods on $x(t)$. For $y(t)$, the contrast is stronger: the ellipsoidal filter achieves the lowest RMSE over the mid- to high-SNR range, while $k$-NN and spherical VR also surpass the time-domain baselines once the SNR is moderate. This behaviour reflects a trade-off: time-domain smoothing reduces noise but also attenuates peak-to-peak structure, so its RMSE gains saturate, whereas topology-based neighbourhoods preserve structure and therefore continue to improve as noise decreases.
@@
Figure 4: Comparison of denoising filters for the two-component signal at SNR = 20 dB. Left: high-amplitude component $x(t)$; right: low-amplitude component $y(t)$. In both panels, (a) original (clean) and noisy signals, (b) moving average, (c) adaptive moving average, (d) $k$-NN topological filter, (e) spherical topological filter, and (f) ellipsoidal topological filter.
Figure 5: RMSE versus SNR for the chirp signal components: top panel shows the high-amplitude component $x(t)$, and bottom panel the low-amplitude component $y(t)$.
V.3 3D accelerometer data denoising
Figure 6: Persistence diagrams computed on the 3D accelerometer signal using Vietoris–Rips (spherical, left) and flow-aware ellipsoidal (right) filtrations. Under the Vietoris–Rips filtration, two $H_{1}$ loops with similar persistence appear close to one another, whereas the ellipsoidal filtration yields a clearer separation of the dominant $H_{1}$ feature from secondary loops. The birth–death interval of the most persistent $H_{1}$ feature is used to define the filtration scale range for the denoising experiments (Figs. 7 and  12).
[…]
@@
We now evaluate the method on real data: publicly available tri-axial accelerometer recordings from cycling (Vieten and Weich, 2020). The data were sampled at 500 Hz; we analysed a 3,500-sample segment from the first participant.
[…]
@@
Fig. 6 shows persistence diagrams for the same 3D segment under the Vietoris-Rips and the proposed Ellipsoidal filtrations (red: $H_{0}$, blue: $H_{1}$). In the VR diagram, two $H_{1}$ classes are clearly more persistent, whereas in the ellipsoidal diagram a single $H_{1}$ class dominates. The blue points farthest from the diagonal represent the most persistent loop in each panel. In the denoising and recurrence time estimation (Section V.4), we consider the scale of the most persistent $H_{1}$ class.
@@
Fig. 7 compares spherical and ellipsoidal neighbourhood averaging across four scales. At small scales, both methods preserve the data’s loop structure. As the scale increases, spherical neighbourhoods span opposite sides of narrow passages and average across bottlenecks, compressing the geometry and attenuating amplitude. In contrast, ellipsoidal averaging follows the local flow and better preserves loops across scales.
@@
Fig. 12 applies denoising by taking the geometric median of the samples within each sphere or ellipsoid. The geometric median yields smoother reconstructions and exhibits greater robustness to noise than the mean.
V.4 Recurrence-time estimation
@@
We assess whether the same flow-aware neighbourhoods that improved denoising (Sections V.2V.3) also improve recurrence-time estimation. The concept of recurrence and first-return times was introduced in Section II.1. For consistency, we use the chirp signal defined in Section V.2, which combines a linearly increasing frequency, anisotropic deformation, and a narrow bottleneck as described earlier.
@@
$aligned T_{1}(i)=\min\{\,j>i:\;&j-i\geq\tau_{\min},\ x_{j}\in\mathcal{N}_{i},\\ &x_{k}\notin\mathcal{N}_{i}\ \text{for }i<k<j\}.$ (21)
[…]
Fig. 8 presents the persistence diagrams of the clean and noisy signals for both Vietoris–Rips and ellipsoidal filtrations, highlighting the dominant $H_{1}$ features under each setting.
@@
Figs. 10a and 10b compare first-return-time estimates across persistence-normalised scales. On the clean signal, ellipsoidal neighbourhoods match the ground truth most closely, and at the $H_{1}$ death scale recover the first return for nearly all indices. In contrast, fixed-radius balls struggle as the period shortens: first returns are frequently missed at all scales, and at larger scales balls span opposite sides of the bottleneck, producing spurious early returns. Under noise, ellipsoidal neighbourhoods capture more first returns. Importantly, no spurious early returns are caused by the bottleneck which makes it possible to estimate returns for all points.
Because the ellipsoidal neighbourhoods are point-specific, the induced recurrence relation is generally asymmetric (i.e. $x_{j}\in\mathcal{N}_{i}$ does not imply $x_{i}\in\mathcal{N}_{j}$), which naturally leads to a directed recurrence plot. In this work, we focus on first-return times and leave the construction of directed recurrence plots, and the corresponding Recurrence Quantification Analysis (RQA) measures, to future work.
Figure 8: Persistent diagrams for clean (top row) and noisy (bottom row) signals under Vietoris–Rips filtration (left column) and flow-aware ellipsoidal filtration (right column). Noise introduces additional low-persistence $H_{1}$ features, visible as points closer to the diagonal. In each diagram, the $H_{1}$ point farthest from the diagonal corresponds to the most persistent loop. Its birth–death interval defines the dominant $H_{1}$ persistence range used to select the filtration scales for recurrence-time estimation in Fig. 10a.
a
b
[…]
VI Discussion